OpenStudy (anonymous):

solve for x. x/5 = 1 1/2

OpenStudy (imtiaz7):

x/5= 11/2 2x=55 x=55/2

OpenStudy (anonymous):

okay, thank you!

geerky42 (geerky42):

\(1\dfrac{1}{2}\neq\dfrac{11}{2}\)

OpenStudy (imtiaz7):

@denise_sega you are forgeting something

geerky42 (geerky42):

lol medal grabber.

OpenStudy (imtiaz7):

lol

geerky42 (geerky42):

If right side is meant to be \(1\dfrac{1}{2}\), then just convert it to improper fraction, which is \(\dfrac{3}{2}\) So in that case, just multiply both sides by 5 and you will have \(\boxed{x=\dfrac{15}{2}}\)